Solve the differential equation:
xdy - (3y + x^5. y^1/3)dx = 0
1
Expert's answer
2020-03-06T15:56:56-0500

"xdy-(3y+x^5y^{\frac{1}{3}})dx=0\\
xy'-3y=x^5y^{\frac{1}{3}}"

Bernoulli equation

divisible by "y^{\frac{1}{3}}"

"\frac{xy'}{y^{\frac{1}{3}}}-3y^\frac{2}{3}=x^5"

we will replace

"y^{\frac{2}{3}}=z(x)\\
y=z^{\frac{3}{2}}\\
y'=\frac{3}{2}z^{\frac{1}{2}}z'\\
y^{\frac{1}{3}}=z^{\frac{1}{2}}"

then

"\frac{x\cdot\frac{3}{2}\cdot z^{\frac{1}{2}}\cdot z'}{z^{\frac{1}{2}}}-3z=x^5\\
\frac{3}{2}xz'-3z=x^5\\
3xz'-6z=2x^5"

obtain a linear equation

1)

"3xz'-6z=0\\
\frac{dz}{z}=\frac{2dx}{x}\\
ln|z|=2ln|x|+ln|C|\\
z=C\cdot x^2"

2)

"z=C(x)\cdot x^2\\
z'=C'x^2+2Cx\\"

substitute into the equation

"3xz'-6z=2x^5"

"3x(C'x^2+2Cx)-6Cx^2=2x^5\\
3x^3C'-6x^2C-6Cx^2=2x^5\\
C'=\frac{2}{3}x^2\\
C(x)=\frac{2}{3}\cdot\frac{x^3}{3}+C_1"

Then

"z=(\frac{2x^3}{9}+C_1)x^2\\
y^{\frac{2}{3}}=(\frac{2x^3}{9}+C_1)x^2\\" is the general solution,

 "y=0" is a solution